OpenStudy (anonymous):

In a survey of 500 people, 200 indicated that they would be buying a major appliance within the next month. 150 indicated that they would buy a car, and 25 said that they would purchase both a major appliance and a car. How many will purchase neither. How many will purchase only a car?

OpenStudy (kropot72):

Number buying only a car = 150 - 25 = 125 Number buying only a major appliance = 200 - 25 = 175 Number making a purchase of either a car or major appliance or both = 125 + 175 + 25 = 325 Number purchasing neith a car or a major appliance = 500 - 325 = 175